OpenStudy (anonymous):

For what integer k, k>1, will both sigma infinity when n=1 ((-1)^(kn))/n and sigma infinity when n=1 (k/4)^n converge?

OpenStudy (anonymous):

Have you tried the alternating series test for the first series?

OpenStudy (anonymous):

\[\large\sum_{n=1}^\infty\frac{(-1)^{kn}}{n}\text{ will converge if}\\ \lim_{n\to\infty}\frac{1}{n}=0, \text{ and}\\ \left\{\frac{1}{n}\right\}\text{ is a decreasing sequence.}\] However, this series will only converge if k is odd (1, 3, 5, ...). Otherwise, you have the series \[\sum_{n=1}^\infty\frac{1}{n}, \text{ which diverges.}\]

OpenStudy (anonymous):

The second series is a geometric series. What do you know about the convergence conditions of a geo-series?